OpenStudy (anonymous):

2√10 times 3√12 radical expression?

OpenStudy (cwrw238):

2*3 = 6 and √10 * √12 = √120 = 6 √120 now you can simplify √120 by finding its prime factors and checking for perfect squares

OpenStudy (anonymous):

Do i have to simplify the 6 too?

OpenStudy (cwrw238):

nope

OpenStudy (anonymous):

Oh okay

OpenStudy (cwrw238):

120 = 2*2*2*3*5

OpenStudy (cwrw238):

can you make a square number out of these factors?

OpenStudy (anonymous):

i only have one answer that have 6 on the outside of the square root which is 6√120

OpenStudy (anonymous):

no by the way

OpenStudy (anonymous):

and the others are 12√60 5√120 and 12√30

OpenStudy (cwrw238):

ok the square number in the factors is 2*2 = 4 and √4 is 2 so we can write √120 = √4 √30 = 2√30 finally the simplified form is 6 * 2 √30 = 12√30

OpenStudy (cwrw238):

do you follow what I have written ?

OpenStudy (anonymous):

Thank You So Much ! :T And Yes

OpenStudy (cwrw238):

yw